Find the values of a and b that make f continuous everywhere,


f(x)={((x^(2)+6x+5)/(x^(2)-3x-4),x<-1),(ax^(2)-bx+3,-1<=x<3),(2x-a+b,x>=3):}

1
Expert's answer
2021-11-17T18:18:56-0500

f (x)={(( x2+6x+5)/(x2-3x-4),x <-1), ax2-bx+3,-1 <x <3),(2x-a+b,x>=3)}


x2+ 6x+ 5= x2+ x+ 5x+ 5

=x (x+1) + 5 (x+1)

= (x+5) (x+1)


x2- 3x- 4= x2- 4x+ x- 4

x (x-4)+1 (x-4)= (x+1) (x-4)


Therefore, (x+5) (x+1)/ (x+1) (x-4)

=(x+5)/ (x-4)

x cannot be equal to 4 since the denominator will equate to zero.

Taking x= -1; f (x)=(-1+ 5) / (-1- 4)

=-4/ 5

limx~-1 ax2- bx+ 3= -4/ 5

a+ b=-19/5, let this be equation 1.


limx~3 ax2-bx+3=-4/5

9a-3b=-19/5, let this be equation 2.


limx~3 2x- a+ b= -4/5

-a+ b= -34/5, let this be equation 3.


Taking equation 1 and 2;

9a-3b= -19/5

a+b= -19/5


9a-3b= -19/5

9a+27b= -171/5; substracting the 2 equations;

-30b= 152/5

b= 76/75


a+b=-19/5

a+76/75=-19/5


a+76/75=-19/5

a=-361/75


a=-361/75

b=76/75